OpenStudy (tgk):

Indicate in standard form the equation of the line passing through the given points. A(4, 1), B(5, 2) it has to fit on these lines ___=______

OpenStudy (tgk):

i just really need help trying to graduate

OpenStudy (welshfella):

first find the slope of the line can you do that?

OpenStudy (welshfella):

slope = (y2 - y1) / (x2 - x1) where the 2 points are given by (x1,y1) and (x2,y2)

OpenStudy (tgk):

so would it be: (2-5) divided by (1-4)= -7?

OpenStudy (hima):

IF A(4, 1), B(5, 2) then let A(x1, y1) and B(x2,y2) where slope= (y2 - y1) / (x2 - x1) AS welshfella said then slope = (2-1) / (5-4) = 1/1=1 the next step get the equation form, work on that